# The IACOB project

### v1

@@ 6 Conclusions and future work
In this work, we considered 670 spectroscopically-observed Galactic O-stars and BSGs. Using the ALS III catalog as reference, we found that our sample comprises 60% of all massive stars within 2500 pc and ¡ 11 that are born as O-type stars and reach the spectral type B6. This results in the most complete sample of Galactic massive stars spectroscopically analyzed to date. Combining the results of the quantitative spectroscopic analysis with Gaia distances, we were able to derive their absolute luminosities. This allowed us to represent the sample in the HR diagram and to analyze its properties in the context of stellar evolution.
@@
Based on the distribution of values with , we argue that the evolution and the larger fraction of the slowly-rotating component may be an indicator that massive stars enter the MS with low to mild initial rotations (/ 0.2). This would be further supported under the assumption that the fast-rotating component mainly emerges from binary interaction.
@@
The results presented in this work provide strong empirical constraints for improving stellar evolution models. Particularly with respect to the width of the MS in the massive star domain. Although our results also agree with the possibility that fast-rotating stars may have binary origin, further investigation is required to confirm this hypothesis. Furthermore, stars evolving blueward may also populate the upper part of the HR diagram near the TAMS. Determining the surface abundances and analyzing the pulsation properties of the stars in our sample will help us to find new clues on their evolutionary nature, and will be addressed in a future work. Moreover, the fraction of spectroscopic binaries within our sample demands further investigation and will likely add important constraints on the frequency of binary interaction in massive star evolution.
@@ Acknowledgements.
The authors thank the anonymous reviewer for a constructive and positive report that has led to the improvement of the manuscript. The authors also thank Dominic Bowman for providing valuable comments on this manuscript. AdB, SS-D, and GH acknowledge support from the Spanish Ministry of Science and Innovation and Universities (MICIU) through the Spanish State Research Agency (AEI) through grants PID2021-122397NB-C21, PID2022-136640NB-C22, 10.13039/501100011033, and the Severo Ochoa Program 2020-2023 (CEX2019-000920-S). Regarding the observing facilities, this research is based on observations made with the Mercator Telescope, operated by the Flemish Community at the Observatorio del Roque de los Muchachos (La Palma, Spain), of the Instituto de Astrofísica de Canarias. In particular, obtained with the HERMES spectrograph, which is supported by the Research Foundation - Flanders (FWO), Belgium, the Research Council of KU Leuven, Belgium, the Fonds National de la Recherche Scientifique (F.R.S.-FNRS), Belgium, the Royal Observatory of Belgium, the Observatoire de Genève, Switzerland, and the Thüringer Landessternwarte Tautenburg, Germany. This work is also based on observations with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University, at the Observatorio del Roque de los Muchachos, of the Instituto de Astrofísica de Canarias. Additionally this work is based on observations obtained with the FEROS spectrograph attached to the 2.2 m MPG/ESO telescope at the La Silla observatory (Chile). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement.References
@@
- Gaia Collaboration et al. (2022) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2022, arXiv e-prints, arXiv:2208.00211
@@ Appendix A Further notes on the completeness of the sample
In Sect. 3 we evaluated the completeness level of the stars analyzed with respect to the ALS III catalog. Figure 8 extends this by showing the observed and missing stars within 2500 pc, separated by their spectral type. We do this to further evaluate the potential existence of biases in the distribution of stars shown in this work that may lead to misinterpretations of the results.
@@ Appendix B Evidence of disks near critical velocity
Appendix C The minimum observed spin-rates increasing with luminosity
@@ Appendix D Peak-to-peak detection thresholds for SB1 systems
Appendix E Spin-rate properties of stars with = 4.65 – 4.3
The bottom panel of Fig. 4 has some characteristics that are different from the other panels. It is the only panel in which the median values are above the threshold limit of 100 km s-1, resulting from the smaller fraction of stars in the low- component (see the corresponding panels of Fig. 3). Although this characteristic is not new (see, e.g., Hunter et al. 2008; Holgado et al. 2022), the higher number of fast-rotating stars allows us to better see the “diagonal” decrease of in the 25 – 17 kK range. This panel also shows that most of the stars beyond the TAMS are fast-rotating objects. However, we note the lack of slowly-rotating stars in the 21 – 15 kK range, for which we do not find a specific reason, but it could be related to the selection of stars with spectroscopic luminosities with ¿ 3.5 dex in de Burgos et al. (2024b). The inclusion of additional stars in that range would shift the location of the TAMS toward cooler temperatures, resulting in a lesser number of fast-rotating objects beyond the TAMS.

### v2

@@ 6 Conclusions and future work
In this work, we considered 670 spectroscopically observed Galactic O-stars and BSGs. Using the ALS III catalog as reference, we found that our sample comprises 60% of all massive stars within 2500 pc and ¡ 11 that are born as O-type stars and reach the spectral type B6. This results in the most complete sample of Galactic massive stars spectroscopically analyzed to date. Combining the results of the quantitative spectroscopic analysis with Gaia distances, we were able to derive their absolute luminosities. This allowed us to represent the sample in the HR diagram and to analyze its properties in the context of stellar evolution.
@@
Based on the distribution of values with , we argue that the evolution and the larger fraction of the slowly rotating component may be an indicator that massive stars enter the MS with low to mild initial rotations (/ 0.2). This would be further supported under the assumption that the fast-rotating component mainly emerges from binary interaction.
@@
The results presented in this work provide strong empirical constraints for improving stellar evolution models. Particularly with respect to the width of the MS in the massive star domain. Although our results also agree with the possibility that fast-rotating stars may have binary origin, further investigation is required to confirm this hypothesis. Furthermore, stars evolving blueward may also populate the upper part of the HR diagram near the TAMS. Determining the surface abundances and analyzing the pulsation properties of the stars in our sample will help to find new clues regarding their evolutionary nature and will be addressed in a future work. Moreover, the fraction of spectroscopic binaries within our sample demands further investigation and will likely add important constraints on the frequency of binary interaction in massive star evolution.
@@ Acknowledgements.
The authors thank the anonymous reviewer for a constructive and positive report that has led to the improvement of the manuscript. The authors also thank Dominic Bowman for providing valuable comments on this manuscript. AdB, SS-D, and GH acknowledge support from the Spanish Ministry of Science and Innovation and Universities (MICIU) through the Spanish State Research Agency (AEI) through grants PID2021-122397NB-C21, PID2022-136640NB-C22, 10.13039/501100011033, and the Severo Ochoa Program 2020-2023 (CEX2019-000920-S). Regarding the observing facilities, this research is based on observations made with the Mercator Telescope, operated by the Flemish Community at the Observatorio del Roque de los Muchachos (La Palma, Spain), of the Instituto de Astrofísica de Canarias. In particular, obtained with the HERMES spectrograph, which is supported by the Research Foundation - Flanders (FWO), Belgium, the Research Council of KU Leuven, Belgium, the Fonds National de la Recherche Scientifique (F.R.S.-FNRS), Belgium, the Royal Observatory of Belgium, the Observatoire de Genève, Switzerland, and the Thüringer Landessternwarte Tautenburg, Germany. This work is also based on observations with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University, at the Observatorio del Roque de los Muchachos, of the Instituto de Astrofísica de Canarias. Additionally this work is based on observations obtained with the FEROS spectrograph attached to the 2.2 m MPG/ESO telescope at the La Silla observatory (Chile). This work has used data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement.References
@@
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
@@ Appendix A Further notes on the completeness of the sample
In Sect. 3 we evaluated the completeness level of the stars analyzed with respect to the ALS III catalog. Figure 8 extends this by showing the observed and missing stars within 2500 pc, separated by their spectral type. We do this to further evaluate the potential existence of biases in the distribution of the stars shown in this work that may lead to misinterpretations of the results.
@@ Appendix B Evidence of disks near critical velocity
Appendix C The minimum observed spin rates increasing with luminosity
@@ Appendix D Peak-to-peak detection thresholds for SB1 systems
Appendix E Spin rate properties of stars with = 4.65 – 4.3
The bottom panel of Fig. 4 has some characteristics that are different from the other panels. It is the only panel in which the median values are above the threshold limit of 100 km s-1, resulting from the smaller fraction of stars in the low- component (see the corresponding panels of Fig. 3). Although this characteristic is not new (see, e.g., Hunter et al. 2008; Holgado et al. 2022), the higher number of fast-rotating stars allows us to better see the “diagonal” decrease of in the 25 – 17 kK range. This panel also shows that most of the stars beyond the TAMS are fast-rotating objects. However, we note the lack of slowly rotating stars in the 21 – 15 kK range, for which we do not find a specific reason, but it could be related to the selection of stars with spectroscopic luminosities with ¿ 3.5 dex in de Burgos et al. (2024b). The inclusion of additional stars in that range would shift the location of the TAMS toward cooler temperatures, resulting in a lesser number of fast-rotating objects beyond the TAMS.